Break-Even Analysis Annasam Company is a producer of kayaks. There are two lines of kayaks: one for professional outfitters aI need help on the yellow highlighted area.  
Answer #1
Internal rate of return,IRR is the discount rate at which the net present value of both the cash inflows & outflows exactly equals 0(zero)
so, we will find the annual cash flows are:
Year 0 Initial cash outflow equipment purchase) -500000
Yrs. 1-6 After-tax incremental income-- cash inflows (100000*(1-35%)) 65000
Yrs.1-6 Depreciation tax shields(500000*12%*35% 21000
Total 86000
Yr. 6 After-tax salvage cash inflow--(((500000-360000)-50000)*35%)+50000 81500
Now formulating for the IRR equation, where NPV is 0
-500000+(86000*(1-(1+IRR)^-6)/IRR)+(81500/(1+IRR)^6)=0
solving for IRR, we get IRR=
4.85%
As IRR, 4.85% < WACC 7.5%, the project is not recommended.
Cost 500000
Acc. Depn.(500000*12%*6) 360000
carrying value(500000-267798) 140000
Salvage value at end yr.6 50000
Loss on salvage 90000
Tax savings due to loss(90000*35%) 31500
ATCF on salvage(50000+31500) 81500
Considering depn. on declining balance
Year 0 1 2 3 4 5 6
Cost of equipment -500000
AT incl.income 65000 65000 65000 65000 65000 65000
Depn.tax shiled 21000 18480 16262 14311 12594 11082
ATCF on salvage 113771
Total CFs -500000 86000 83480 81262 79311 77594 189853
IRR=
-500000+(86000/(1+IRR)^1)+(83480/(1+IRR)^2)+(81262/(1+IRR)^3)+(79311/(1+IRR)^4)+(77594/(1+IRR)^5)+(189853/(1+IRR)^6)=0
IRR=4.75%
If depn. Is on declining balance
Year Carrying value Depn. Tax shield=Depn.*tax Rate
1 500000 60000 21000
2 440000 52800 18480
3 387200 46464 16262
4 340736 40888 14311
5 299848 35982 12594
6 263866 31664 11082
Acc. Depn. 267798 93729
carrying value(500000-267798) 232202
Salvage value at end yr.6 50000
Loss on salvage 182202
Tax savings due to loss(182202*35%) 63771
ATCF on salvage(50000+63771) 113771